6.  Problem 8.08 (Beta Coefficient)

eBook

Given the following information, determine the beta coefficient for Stock L that is consistent with equilibrium: r̂L = 9.5%; rRF = 1.5%; rM = 13.5%. Round your answer to two decimal places.

7.  Problem 8.09 (Required Rate of Return)

Stock R has a beta of 1.5, Stock S has a beta of 0.55, the required return on an average stock is 13%, and the risk-free rate of return is 6%. By how much does the required return on the riskier stock exceed the required return on the less risky stock? Round your answer to two decimal places.

  %
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

6.

9.5=1.5+Beta*(13.5-1.5)

(9.5-1.5)=Beta*(13.5-1.5)

Beta=(9.5-1.5)/(13.5-1.5)

=0.67(Approx)

7.

Required return for:

R=6+1.5*(13-6)=16.5%

S=6+0.55*(13-6)=9.85%

Hence excess Required return=16.5-9.85

=6.65%